Consider the chemical reaction 2NH3(g) ⇌ N2(g) + 3H2(g). The equilibrium is to be established in a 1.0 L container at 1,000 K, where Kc = 4.0 × 10–2. Initially, 1,220 moles of NH3(g) are present. Estimate the equilibrium concentration of N2(g). 


1
Expert's answer
2022-03-13T15:21:29-0400

Solution:

Molarity of NH3 = Moles of NH3 / Volume of container = (1220 mol) / (1.0 L) = 1220 M


Balanced chemical equation:

2NH3(g) ⇌ N2(g) + 3H2(g)

From this, we can construct an ICE table:



The Kc expression for the reaction is:



From the ICE Table:

[NH3] = 1220 − 2X ≈ 1220 (since 1220 >> 2X)

[N2] = X

[H2] = 3X

Kc = 4.0×10–2 (according to the task)


Therefore,




(4.0×10–2) × (1488400) = 27X4

59536 = 27X4

X4 = 2205.04

X = (2205.04)1/4 = 6.85


Therefore,

[NH3] = (1220 − 2X) M = (1120 − 2×6.85) M = 1206.3 M

[H2] = (3X) M = (3×6.85) M = 20.55 M

[N2] = X = 6.85 M


Answer: The equilibrium concentration of N2(g) is 6.85 M